OpenStudy (anonymous):

6-5/4x=-4 solve and show work

OpenStudy (anonymous):

-16x=1 x=-1/16

OpenStudy (anonymous):

is that answer?

OpenStudy (anonymous):

6-5=-16x(by cross multiplying) 1=-16x x=-1/16